# Faster Quantum Simulation Of Markovian Open Quantum Systems Via Randomisation

### v2

@@ 6.1.3 Implementation of QDRIFT Channel with CS
Figure 4: Quantum circuit depicting the implementation of the QDRIFT channel using CS. Here each $j_{k}$ is sampled from the discrete distribution $p_{k}$ as defined in equation (112)
@@
2:$N\leftarrow\left\lceil(t\Gamma\Omega)^{2}/\epsilon\ \right\rceil$
3:$\omega\leftarrow t\Gamma/N$
4:$l\leftarrow 0$
5:$G_{QD}=\{\}$
@@
Consider the QDRFIT channel $\mathcal{E}_{\omega}^{(QD)}$ as in (113). By definition, it will apply $\exp(\omega\mathcal{L}_{k})$ with some probability $p_{k}$ with $k=1,2,...,M$. This gives us an easy way to construct a quantum circuit that implements $(\mathcal{E}_{\omega}^{(QD)})^{N}$ using classical sampling. We start by defining a random variable $j_{l}\in\{1,...,M\}$, with $l=1,...,N$, and the probability distribution $p(j_{l}=k)=p_{k}$ with $k=1,...,M$. Then we denote the gate set by $G_{QD}$ and define the classical oracle function SAMPLE() which samples from the distribution $p(j_{l}=k)=p_{k}$ for $k=1,...,M$ and returns a value from the set $\{1,...,M\}$. To construct the circuit we iteratively use SAMPLE() to find a value $j_{l}$ and append $\exp(\omega\mathcal{L}_{j_{l}})$ for $l=1,...,N$ to the gate set $G_{QD}$. Once we have constructed the gate set, we apply it to the initial sate $\rho(0)$, to approximate $\rho(t)$ to a precision $\epsilon=(t\Gamma\Omega)^{2}/N$. Algorithm 3 outlines the pseudocode for constructing the gate set $G_{QD}$ to implement $(\mathcal{E}_{\omega}^{(QD)})^{N}$. Figure 4. shows the quantum circuit that implements the QDRIFT channel, we see here that each exponential depends on the outcome of sampling from the distribution $p_{k}$. Since the $G_{QD}$ contains only exponentials the gate complexity just scales here with the number of elements in $G_{QD}$ i.e. $O((t\Gamma\Omega)^{2}/\epsilon)$.
@@ 6.2.1 Quantum Circuit Implementation of $S_{1}^{(ran)}$ with QF
In this section, we present a method for implementing the randomised TS formula $S_{1}^{(ran)}(\tau)^{N}$ on a quantum computer without the need for classical sampling. We will make use of the quantum forking procedure [30] to construct a quantum circuit that directly implements $S_{1}^{(ran)}(\tau)$, which can be seen in Figure 5. The circuit makes use of controlled swap channels denoted by the usual circuit notation for the controlled-SWAP operation as seen in Figure 5. The controlled swap channel will only swap the states if the state that it is controlled on is in the state |$\outerproduct{1}{1}$ |. We perform $N$ repetitions of the circuit in Figure 5. to obtain $S_{1}^{(ran)}(\tau)^{N}$. The following lemma shall show how the circuit in Figure 5 implements $S_{1}^{(ran)}(\tau)$ using quantum forking.
@@
$\rho_{\text{prep}}\otimes\rho(0)\otimes\rho_{\phi}=\frac{1}{2}\left(\outerproduct{0}{0}\otimes\rho(0)\otimes\rho_{\phi}+\outerproduct{1}{1}\otimes\rho(0)\otimes\rho_{\phi}\right),$ (132)
@@
$\frac{1}{2}\left(\outerproduct{0}{0}\otimes\rho(0)\otimes\rho_{\phi}+\outerproduct{1}{1}\otimes\rho_{\phi}\otimes\rho(0)\right).$ (133)
@@
$\frac{1}{2}\left(\outerproduct{0}{0}\otimes S_{1}^{\rightarrow}(\tau)\rho(0)\otimes S_{1}^{\leftarrow}(\tau)\rho_{\phi}+\outerproduct{1}{1}\otimes S_{1}^{\rightarrow}(\tau)\rho_{\phi}\otimes S_{1}^{\leftarrow}(\tau)\rho(0)\right),$ (134)
@@
$\frac{1}{2}\left(\outerproduct{0}{0}\otimes S_{1}^{\rightarrow}(\tau)\rho(0)\otimes S_{1}^{\leftarrow}(\tau)\rho_{\phi}+\outerproduct{1}{1}\otimes S_{1}^{\leftarrow}(\tau)\rho(0)\otimes S_{1}^{\rightarrow}(\tau)\rho_{\phi}\right).$ (135)
@@
$\frac{1}{2}\left(\mathrm{tr}(\outerproduct{0}{0})\otimes S_{1}^{\rightarrow}(\tau)\rho(0)\otimes\mathrm{tr}(S_{1}^{\leftarrow}(\tau)\rho_{\phi})+\mathrm{tr}(\outerproduct{1}{1})\otimes S_{1}^{\leftarrow}(\tau)\rho(0)\otimes\mathrm{tr}(S_{1}^{\rightarrow}(\tau)\rho_{\phi})\right).$ (136)
@@
Now that we understand how to implement $S_{1}^{(ran)}$ for a small time step using quantum forking, we need to show that if we repeat this circuit $N=\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ times we will implement $S_{1}^{(ran)}(\tau)^{N}$. The circuit in Figure 6 illustrates how one can repeat the circuit in Figure 5. It relies on two important operations. The first important operation is the measurement operation with discard correspond to tracing out the respective register, and the second important operation is represented by the grey bar with a dotted line, this is called a barrier. It represents the process of resetting the register to a desired state after measuring and discarding the result. The following theorem will outline how the circuit in Figure 6 implements $S_{1}^{(ran)}(\tau)^{N}$.
@@
Repeating this process above $N-2$ times we have that the final state of the system register is $S_{1}^{(ran)}(\tau)^{N}\rho(0)$, which is the desired output of the circuit. Now if we define $\tilde{\rho}(t)=S_{1}^{(ran)}(\tau)^{N}\rho(0)$ and make use of Theorem 3 we have
@@
We can now determine the gate complexity for the quantum circuit that implements $S_{1}^{(ran)}(t/N)^{N}$. Unlike the usual definition of gate complexity where one counts the number of elementary gates used from a universal set, we have chosen to count the number of simple quantum channels implemented in the circuit. For example in the circuit in Figure 5, the channel $S_{1}^{\rightarrow}(t/N)$ contains $M$ exponentials and there are $M$ exponentials in $S_{1}^{(\leftarrow)}(t/N)$. There are also two controlled-SWAP channels in the circuit which we will need to count as well. Therefore to implement $S_{1}^{(ran)}(t/N)$, we will require $22M+$ 2 simple channels. If we need to repeat this circuit $N=\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ times then in total we will need $(2M+2)\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ number of simple channels to implement $S_{1}^{(ran)}(t/N)^{N}$. If we define the gate complexity for the first order randomised TS product formula as $g_{1}^{(ran)}$ then the gate complexity scales as
@@ 6.2.3 Quantum Circuit Implementation of QDRIFT Channel with QF
$\rho_{\text{prep}}\otimes\rho(0)\otimes(\rho_{\phi})^{\otimes M-1}=\sum_{k=1}^{M}p_{k}(\outerproduct{k}{k}\otimes\rho(0)\otimes\rho_{\phi}\otimes...\otimes\rho_{\phi}),$ (142)
where the states $\rho_{\phi}$ are once again just easy to prepare states that have no effect on the outcome of the circuit. The controlled-SWAP channels in Figure 8 are applied only when the ancillary register $\rho_{\text{prep}}$ is in the state |$\ket{k}\bra{k}$ |. Applying all the controlled-SWAP channels produces the state
$p_{1}(\outerproduct{1}{1}\otimes\rho(0)\otimes\rho_{\phi}\otimes...\otimes\rho_{\phi})+p_{2}(\outerproduct{2}{2}\otimes\rho_{\phi}\otimes\rho(0)\otimes...\otimes\rho_{\phi})+...$
$...+p_{M-1}(\outerproduct{M-1}{M-1}\otimes\rho_{\phi}\otimes...\otimes\rho(0)\otimes\rho_{\phi})+p_{M}(\outerproduct{M}{M}\otimes\rho_{\phi}\otimes\rho_{\phi}\otimes...\otimes\rho(0)).$ (143)
@@
$p_{1}(\outerproduct{1}{1}\otimes e^{\omega\mathcal{L}_{1}}\rho(0)\otimes e^{\omega\mathcal{L}_{2}}\rho_{\phi}\otimes...\otimes e^{\omega\mathcal{L}_{M}}\rho_{\phi})+p_{2}(\outerproduct{2}{2}\otimes e^{\omega\mathcal{L}_{1}}\rho_{\phi}\otimes e^{\omega\mathcal{L}_{2}}\rho(0)\otimes...\otimes e^{\omega\mathcal{L}_{M}}\rho_{\phi})+...$
$...+p_{M}(\outerproduct{M}{M}\otimes e^{\omega\mathcal{L}_{1}}\rho_{\phi}\otimes e^{\omega\mathcal{L}_{2}}\rho_{\phi}\otimes...\otimes e^{\omega\mathcal{L}_{M}}\rho(0)).$ (144)
@@
$p_{1}(\outerproduct{1}{1}\otimes e^{\omega\mathcal{L}_{1}}\rho(0)\otimes e^{\omega\mathcal{L}_{2}}\rho_{\phi}\otimes...\otimes e^{\omega\mathcal{L}_{M}}\rho_{\phi})+p_{2}(\outerproduct{2}{2}\otimes e^{\omega\mathcal{L}_{2}}\rho(0)\otimes e^{\omega\mathcal{L}_{1}}\rho_{\phi}\otimes...\otimes e^{\omega\mathcal{L}_{M}}\rho_{\phi})+...$
$...+p_{M}(\outerproduct{M}{M}\otimes e^{\omega\mathcal{L}_{M}}\rho(0)\otimes e^{\omega\mathcal{L}_{2}}\rho_{\phi}\otimes...\otimes e^{\omega\mathcal{L}_{1}}\rho_{\phi}).$ (145)
By tracing out the ancillary registers, we see that $\mathrm{tr}(\outerproduct{k}{k})=\innerproduct{k}{k}=1$. Then since
$\mathrm{tr}(\exp(\omega\mathcal{L}_{k})\rho_{\phi})=1$ for any $k=1,...,M$ we have the output state
@@
We are now able to find the gate complexity of the circuit that implements the QDRIFT channel. If we consider the circuit in Figure 7 we see that to implement $\mathcal{E}_{\omega}^{(QD)}$ we need $2(M-1)$ controlled-SWAP channels and we implement $M$ simpler channels i.e. $\exp(\omega\mathcal{L}_{k})$ for $k=1,...,M$. This means that in total to implement $\mathcal{E}_{\omega}^{(QD)}$ we need $23M-$ 2 simple channels. Now if we implement $\mathcal{E}_{\omega}^{(QD)}$, $N=\lceil(t\Gamma\Omega)^{2}/\epsilon\rceil$ times then in total to implement $(\mathcal{E}_{\omega}^{(QD)})^{N}$ we need $(3M-2)\lceil(t\Gamma\Omega)^{2}/\epsilon\rceil$ simple channels. If we denote the gate complexity for the QDRIFT channel as $g^{(QD)}$, then the gate complexity scales as
@@
7 Comparison Of Gate Complexities
In this section we will discuss and compare the gate complexities, for all the simulation methods outline in this paper to the gate complexities of the deterministic TS product formulas. Table 1 summarises the gate complexities for each method.
Method Of Simulation Gate Complexity
Second Order TS Deterministic $O\left((t\Lambda)^{3/2}M^{5/2}/\sqrt{3\epsilon}\right)$
First Order TS Randomised (CS) $O\left((t\Lambda)^{3/2}M^{5/2}/\sqrt{3\epsilon}\right)$
QDRIFT (CS) $O\left((t\Gamma\Omega)^{2}/\epsilon\right)$
First Order TS Randomised (QF) $O\left((t\Lambda)^{3/2}M^{5/2}/\sqrt{3\epsilon}\right)$
QDRIFT (QF) $O\left((t\Gamma\Omega)^{2}M/\epsilon\right)$
As mentioned earlier, the goal in using randomisation to simulate Markovian OQS is to achieve faster simulation, by improving the gate complexity dependence on the number of terms in the generator $\mathcal{L}$, denoted by $M$. In Table 1 we see that the randomised formulas (both CS and QF) and the QDRIFT channel all improve on the gate complexities dependence on $M$ when compared to the deterministic TS product formulas. The first order randomised TS formula (CS) scales the same as the second order deterministic formula. This is an improvement over the first order deterministic TS formula as we see an improvement in the scaling of the gate complexity with respect to the number of terms in the generator $M$, as it now scales as $M^{5/2}$ instead of $M^{3}$. We also observe that we achieve the same scaling when implementing the first order randomised formula with quantum forking. This means that there is no additional cost to performing the sampling directly on the quantum computer. The second order randomised TS (CS) gives us a quadratic dependence on $M$ which is much better than both the deterministic TS product formulas and even the first order randomised TS formula. However, as discussed in section 6.2.2 we cannot use QF to implement $S_{2}^{(ran)}$ due to the fact that we will require $2(M!)$ controlled-SWAP channels which implies that the gate complexity will depend on $M!$ which is very inefficient. For the QDRIFT channel (CS), we observe that there is no dependence on $M$ in the gate complexity. This makes this method ideal for systems with a large number of terms $M$. For example, two and three dimensional Jaynes-Cummings models with neighbour-neighbour interactions [27] and two dimensional Heisenberg models with boundary driving and local dissipation [35]. One may argue however that the $M$ dependence is hidden in the factor $\Gamma$, however for most systems we will compute $\Gamma$ classically before we construct the gate set $G_{QD}$. Therefore, this would not add a factor of $M$ to the gate complexity. For the implementation of the QDRIFT channel with QF we see that we require $2(M-1)$ controlled-SWAP channels which means that the gate complexity will now scale linearly with $M$ which is still better than all other TS product formulas (deterministic and randomised). However, while this scaling with $M$ is better than other methods, the quadratic scaling in $t$ is what makes this method only applicable for short simulation times, for longer simulation times, a randomised product formula will perform better.
@@ 8 Conclusion
An open problem that can be addressed in future is to find optimal convex mixtures of second order TS product formulas that can improve precision and or acheive better scaling of the gate complexity in terms of $M$. A second open problem will be to see if one can apply these results to the simulation of quantum channels that describe non-Markovian dynamics of an OQS [36, 37, 38].

### v3

@@ 6.1.3 Implementation of QDRIFT Channel with CS
Figure 4: Quantum circuit depicting the implementation of the QDRIFT channel using CS. Here each $j_{k}$ is sampled from the discrete distribution $p_{k}$ as defined in equation (107).
@@
[…]
@@
Consider the QDRFIT channel $\mathcal{E}_{\omega}^{(QD)}$ as in (108). By definition, it will apply $\exp(\omega\mathcal{L}_{k})$ with some probability $p_{k}$ with k=1,2,,Mk=1,2,...,M. This gives us an easy way to construct a quantum circuit that implements $(\mathcal{E}_{\omega}^{(QD)})^{N}$ using classical sampling. We start by defining a random variable $j_{l}\in\{1,...,M\}$, with l=1,,Nl=1,...,N, and the probability distribution $p(j_{l}=k)=p_{k}$ with k=1,,Mk=1,...,M. Then we denote the gate set by $G_{QD}$ and define the classical oracle function SAMPLE() which samples from the distribution $p(j_{l}=k)=p_{k}$ for k=1,,Mk=1,...,M and returns a value from the set $\{1,...,M\}$. To construct the circuit we iteratively use SAMPLE() to find a value $j_{l}$ and append $\exp(\omega\mathcal{L}_{j_{l}})$ for l=1,,Nl=1,...,N to the gate set $G_{QD}$. Once we have constructed the gate set, we apply it to the initial sate $\rho(0)$, to approximate $\rho(t)$ to a precision $\epsilon=(t\Gamma\Omega)^{2}/N$. Algorithm 3 outlines the pseudocode for constructing the gate set $G_{QD}$ to implement $(\mathcal{E}_{\omega}^{(QD)})^{N}$. Figure 4. shows the quantum circuit that implements the QDRIFT channel, we see here that each exponential depends on the outcome of sampling from the distribution $p_{k}$. Since the $G_{QD}$ contains only exponentials the gate complexity just scales here with the number of elements in $G_{QD}$ i.e. $O((t\Gamma\Omega)^{2}/\epsilon)$.
@@ 6.2 Quantum Circuit Implementation of $S_{1}^{(ran)}$ and QDRIFT via Quantum Forking
In this section, we explore Quantum Forking (QF) [32] as a method for implementing the randomised algorithms directly on a quantum computer, thereby removing the need for a classical co-processor to perform the sampling. The primary motivation for this approach is to provide a ”quantum-native” pathway where the randomisation is an intrinsic part of the quantum circuit itself. This is conceptually significant in the study of quantum algorithms, as it avoids classical-quantum communication within the main simulation loop. However, this approach involves a clear trade-off which we now explicitly discuss. While the CS implementation of QDRIFT offers superior gate complexity that is independent of $M$, the QF implementation is presented as a method to perform the randomisation natively on the quantum computer, albeit at the cost of reintroducing a linear dependence on $M$. This cost-benefit analysis is crucial when selecting an implementation strategy.
@@ 6.2.1 Quantum Circuit Implementation of $S_{1}^{(ran)}$ with QF
In this section, we present a method for implementing the randomised TS formula $S_{1}^{(ran)}(\tau)^{N}$ on a quantum computer without the need for classical sampling. We will make use of the quantum forking procedure [32] to construct a quantum circuit that directly implements $S_{1}^{(ran)}(\tau)$, which can be seen in Figure 5. The circuit makes use of controlled swap channels denoted by the usual circuit notation for the controlled-SWAP operation as seen in Figure 5. The controlled swap channel will only swap the states if the state that it is controlled on is in the state $|\outerproduct{1}{1}$. We perform $N$ repetitions of the circuit in Figure 5. to obtain $S_{1}^{(ran)}(\tau)^{N}$. The following lemma shall show how the circuit in Figure 5 implements $S_{1}^{(ran)}(\tau)$ using quantum forking.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Now that we understand how to implement $S_{1}^{(ran)}$ for a small time step using quantum forking, we need to show that if we repeat this circuit $N=\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ times we will implement $S_{1}^{(ran)}(\tau)^{N}$. The circuit in Figure 6 illustrates how one can repeat the circuit in Figure 5. It relies on two important operations. The first important operation is the measurement operation with discard correspond to tracing out the respective register, and the second important operation is represented by the grey bar with a dotted line, this is called a barrier. It represents the process of resetting the register to a desired state after measuring and discarding the result.
This iterative structure relies on mid-circuit measurement and reset operations. We note that it is possible, in principle, to design a circuit with only one measurement block at the final step by applying the principle of deferred measurement. Such an approach would require introducing a new set of ancilla registers for each of the N steps to store the random outcomes. However, the presented method, with its explicit trace-out and reset operations, more clearly illustrates the step-by-step implementation of the randomized protocol and simplifies the corresponding proof of Theorem 6.
The following theorem will outline how the circuit in Figure 6 implements $S_{1}^{(ran)}(\tau)^{N}$.
@@
Repeating this process above N2N-2 times we have that the final state of the system register is $S_{1}^{(ran)}(\tau)^{N}\rho(0)$, which is the desired output of the circuit. Now if we define $\tilde{\rho}(t)=S_{1}^{(ran)}(\tau)^{N}\rho(0)$ and make use of Theorem 3 we have
@@
We can now determine the gate complexity for the quantum circuit that implements $S_{1}^{(ran)}(t/N)^{N}$. Unlike the usual definition of gate complexity where one counts the number of elementary gates used from a universal set, we have chosen to count the number of simple quantum channels implemented in the circuit. For example in the circuit in Figure 5, the channel $S_{1}^{\rightarrow}(t/N)$ contains $M$ exponentials and there are $M$ exponentials in $S_{1}^{(\leftarrow)}(t/N)$. There are also two controlled-SWAP channels in the circuit which we will need to count as well. Therefore to implement $S_{1}^{(ran)}(t/N)$, we will require $2M+2$ simple channels. If we need to repeat this circuit $N=\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ times then in total we will need $(2M+2)\lceil(Mt\Lambda)^{3/2}/\sqrt{3\epsilon}\rceil$ number of simple channels to implement $S_{1}^{(ran)}(t/N)^{N}$. If we define the gate complexity for the first order randomised TS product formula as $g_{1}^{(ran)}$ then the gate complexity scales as
@@ 6.2.3 Quantum Circuit Implementation of QDRIFT Channel with QF
[…]
where the states $\rho_{\phi}$ are once again just easy to prepare states that have no effect on the outcome of the circuit. The controlled-SWAP channels in Figure 8 are applied only when the ancillary register $\rho_{\text{prep}}$ is in the state $|\ket{k}\bra{k}$. Applying all the controlled-SWAP channels produces the state
[…]
@@
[…]
@@
[…]
$\mathrm{tr}(\exp(\omega\mathcal{L}_{k})\rho_{\phi})=1$ for any k=1,,Mk=1,...,M we have the output state
@@
We are now able to find the gate complexity of the circuit that implements the QDRIFT channel. If we consider the circuit in Figure 7 we see that to implement $\mathcal{E}_{\omega}^{(QD)}$ we need 2(M1)2(M-1) controlled-SWAP channels and we implement $M$ simpler channels i.e. $\exp(\omega\mathcal{L}_{k})$ for k=1,,Mk=1,...,M. This means that in total to implement $\mathcal{E}_{\omega}^{(QD)}$ we need 3M23M-2 simple channels. Now if we implement $\mathcal{E}_{\omega}^{(QD)}$, $N=\lceil(t\Gamma\Omega)^{2}/\epsilon\rceil$ times then in total to implement $(\mathcal{E}_{\omega}^{(QD)})^{N}$ we need $(3M-2)\lceil(t\Gamma\Omega)^{2}/\epsilon\rceil$ simple channels. If we denote the gate complexity for the QDRIFT channel as $g^{(QD)}$, then the gate complexity scales as
@@ 7 Conclusion
It is important to contextualize our results within the broader landscape of Lindbladian simulation. As shown in our expanded comparison in Table 1, state-of-the-art deterministic algorithms such as Effective Hamiltonian Methods [28] and Repeated Interactions [27] can achieve superior asymptotic scaling in simulation time $t$, approaching a near-linear dependence. However, these advanced methods often rely on complex primitives like block-encodings or quantum signal processing, which can introduce significant constant-factor overheads in terms of gate counts and ancillary qubits. In the era of early fault-tolerant quantum computing, these overheads may be prohibitive. The randomized methods presented in this work, while having less favourable scaling in $t$, offer a compelling alternative due to their structural simplicity, lower implementation overhead, and direct composition from elementary channel operations. Therefore, they represent a practical and resource-efficient choice for near-term simulations, particularly for systems with a large number of generator terms $M$, where our randomized formulas provide a significant polynomial improvement over their direct deterministic counterparts.
A key technical contribution of this work is the direct derivation of error bounds for these randomised methods, which we achieved without relying on the Campbell and Hastings mixing lemma [30, 31]. This direct approach was necessary because the mixing lemma is specific to the geometry of unitary operators and does not generalize to the CPTP maps that describe open system dynamics. While our Taylor expansion and bounding technique is general enough to be reapplied to the special case of Hamiltonian simulation, it would likely yield a more complex proof and potentially looser bounds than those achieved with the elegant mixing lemma. Given that tight bounds for randomised Hamiltonian simulation are already well-established, reapplying our more general method would not provide a clear benefit in that context. However, an interesting open problem that can be addressed is the generalisation of the mixing lemma for CPTP maps.
An open problem that can be addressed in future is to find optimal convex mixtures of second order TS product formulas that can improve precision and or acheive better scaling of the gate complexity in terms of $M$. Another open problem will be to see if one can apply these results to the simulation of quantum channels that describe non-Markovian dynamics of an OQS [40, 41, 42]. Furthermore, a natural extension would be to develop randomised formulas of an order higher than two, analogous to the methods for Hamiltonian simulation [11]. The primary bottleneck for such an extension is the recursive construction of higher-order Suzuki formulas, which requires propagators with negative time steps. While evolving a Hamiltonian for a negative time corresponds to applying the adjoint unitary, the backward-in-time evolution generated by a Lindbladian, $e^{-t\mathcal{L}}$, is not in general a completely positive and trace-preserving (CPTP) map. This loss of physicality prevents a direct generalisation of higher-order randomised formulas to the open systems setting and overcoming it remains a significant challenge.